If a U.S. resident just sold shares in a British firm that had a 15% return (in pounds) during a period when the pound depreciated 5%, the dollar return of the investment is:

  1. 10%
  2. 15%
  3. 9.25%
  4. 5%
  5. none of the above
Answer #1

Answer: C

The formula is as below:

Dollar return = (1 + local return) (1 – currency depreciated) – 1

                        = (1 + 0.15) (1 – 0.05) – 1

                        = 1.15 × 0.95 – 1

                        = 1.0925 – 1

                        = 0.0925

Now, this is to be multiplied by 100 to get in percentage form.

Dollar return = 0.0925 × 100

                        = 9.25%